x[sup]2[/sup] - 6= 5x+9
1
Expert's answer
2011-01-27T08:30:08-0500
x2 - 6= 5x+9
x2 - 5x - 15= 0
D = 25 + 60 = 75
x1 = (5 + 5√3)/2 = 6.83
x2 = (5 - 5√3)/2 =-1.83